OpenStudy (anonymous):

Solve for x. 4 • 7 – 4 • 10 = 4(x – 10)

OpenStudy (anonymous):

I will give anyone a medal and fan them if they answer the question!

jigglypuff314 (jigglypuff314):

order of operations: PEMDAS so for 4 • 7 – 4 • 10 = 4(x – 10) multiply first 28 - 4 • 10 = 4(x – 10) 28 - 40 = 4(x - 10) then simplify the left side, distribute the right side. add 40 to both sides, divide both sides by 4.

OpenStudy (anonymous):

x=7